# Urysohn width and macroscopic scalar curvature

### v1

@@ 1.1. Background and motivation
To remedy this, Alpert-Balitskiy-Guth [ABG24] propose that one should impose a control on the volume of unit balls in the universal cover of a slightly larger concentric ball. We reproduce their proposal as a definition, with the caveat that it is a slight modification of the usual notion of macroscopic scalar curvature, specifically tailored to investigate Conjecture 1 from a macroscopic point of view.
@@ 1.2. Main results
Certainly, a uniruling is an example of a ruling in this sense. Since $\omega\in H^{2}(M;\mathbf{Z})$, there is a principal circle bundle $\pi:E\to M$ naturally associated to $\omega$, satisfying $c_{1}(E)=\omega$. Furthermore, since $\langle\omega,[\Sigma_{x}]\rangle=\pm 1$, the restriction of $E$ over any ruling sphere is the Hopf fibration. We call $E$ the circle bundle associated to the ruling. Our second key result in this article, used to ensure $E_{n}$ has uniformly positive macroscopic scalar curvature, is the following.
@@
Let $M$ be a closed Riemannian manifold that admits a $3/2$-small ruling. For any $\sigma>0$, there exists a small enough fiber radius $\varepsilon>0$ such that the total space $E$ of the circle bundle $S^{1}({\varepsilon})\to E\to M$ associated to the ruling, equipped with the adapted metric (cf. Definition 2.4) satisfies $\mathrm{mscal}\geq\sigma$.
From Theorem B and Theorem C, we see that the sufficient conditions on the Riemannian manifolds $M_{n}^{d-1}$ needed to establish Theorem A are the following:
@@
    Let $N^{d-1}$ be any $\mathbf{Z}_{2}$-enlargeable $\mathrm{Pin}^{-}$ manifold of dimension $d-1\geq 3$. Then, there exists a cover $N_{n}\to N$ such that $\mathrm{HS}(N_{n};\mathbf{Z}_{2})\geq n$. Choose a uniform small constant $\delta>0$. Let $M_{n}$ be obtained from $N_{n}$ by taking a connect sum with $S^{2}\times S^{d-3}$ at each point of a maximal $\delta$-separated collection of points on $N_{n}$. Then $M_{n}$ admits a (highly singular) small ruling by smoothly embedded $2$-spheres of diameters equal to, upto an error of the order of $\delta$, the diameter of the round $2$-spheres $S^{2}\times\{*\}$ that rule $S^{2}\times S^{d-3}$, which we may choose to be less than $3/2$. The class $\omega\in H^{2}(M_{n};\mathbf{Z})$ may be chosen to be the Poincare dual of the linear combination of the homology classes $\{*\}\times S^{d-3}]$, one from each $S^{2}\times S^{d-3}$ connect summand. For details, see Section 4.1.
@@ Acknowledgments
The authors are grateful to Alexey Balitskiy and Larry Guth for suggesting several ideas and patiently answering their questions at various stages of this work. They are also indebted to Bernhard Hanke and Shihang He for helpful correspondence that led to several key ideas in this article. In particular, He pointed out the article [BOL06] to us, which served as the inspiration for Section 2. They are also grateful to Mike Miller Eismeier for help with the algebraic topology in Section 2. They are also thankful to Dennis Sullivan for sharing his insights on symplectic geometry, some of which are behind Section 1.3.
@@ 2.3. Hopf invariant for circle bundles
By the Wu formula, $\mathrm{Sq}^{2}:H^{i}(M;\mathbf{Z}_{2})\to H^{i+2}(M;\mathbf{Z}_{2})$ is given by $\mathrm{Sq}^{2}(x)=v_{2}(M)\smile x$. Thus, if $M$ admits a $\mathrm{Pin}^{-}$ structure, then $\mathrm{Sq}^{2}=0$.
@@
$\Phi:H^{i}(M;\mathbf{Z}_{2})\to H^{i+2}(\mathrm{Th}(E);\mathbf{Z}_{2}),$
@@ 2.4. Width of circle bundles
The results in the following section are not used in the rest of the article. However, it provides an instructive contrast to the results discussed till now.
@@ 2.5. Urysohn width and Cheeger-Gromov collapsing
We now fix $n\gg C$. Let $g_{E,\varepsilon}=\pi^{*}g_{Z}+\varepsilon^{2}\theta\otimes\theta$ be the adapted metric on $E$ associated to the same auxiliary connection as $g_{E}$, but with fiber radius $\varepsilon<1$. Let $g_{n,\varepsilon}:=p_{n}^{*}g_{E,\varepsilon}$. Then $g_{n,\varepsilon}<{g_{n}}$, therefore $\mathrm{UW}_{3}(E_{n},g_{n,\varepsilon})<\mathrm{UW}(E_{n},g_{n})\leq C$ as well. Thus, even though there is a Cheeger-Gromov collapse $(E_{n},g_{n,\varepsilon})\xrightarrow{\mathrm{GH}}Z_{n}$, we still have
@@ 3. Singular ruling and macroscopic scalar curvature
We record a simple observation inspired by the proof of [BB22, Lemma 3.11]. This observation will not be used per se in the rest of the article, but it shows that the ruling spheres of a small (pre-)ruling on a large manifold must necessarily have complicated patterns of intersection. The rulings constructed in Section 4 affirm this behavior.
Lemma 3.1.
Let $M$ be a Riemannian manifold of dimension $n$ such that $\mathrm{UW}_{n-2}(M)\geq R$. Suppose $M$ admits a $\lambda$-small pre-ruling $\mathfrak{S}$. Then there exists a chain of ruling spheres $\Sigma_{1},\Sigma_{2},\cdots,\Sigma_{N}$ such that $\Sigma_{k}\cap\Sigma_{k+1}\neq\emptyset$ and $N\geq R/\lambda$.
Proof.
Let us define a relation $\approx$ on $M$ by defining $p\approx q$ if there exists ruling spheres $\Sigma_{p},\Sigma_{q}\in\mathfrak{S}$ passing through $p$ and $q$, respectively, such that $\Sigma_{p}$ intersects $\Sigma_{q}$. Let $\sim$ denote the equivalence relation on $M$ given by the transitive closure of $\approx$. Then $M/\sim$ is a simplicial complex of dimension at most n2n-2. Since $\mathrm{UW}_{n-2}(M)>R$, the quotient map $f:M\to M/\sim$ must have a fiber of diameter at least $R$. Since the fibers of $f$ consist of pairwise intersecting ruling spheres, the conclusion follows. ∎
@@
Let $(M^{d},g_{M})$ be a closed Riemannian manifold with a $3/2$-small ruling $\mathfrak{S}$. Let $\pi:E\to M$ be the circle bundle associated to $\mathfrak{S}$. We equip $E$ with an adapted metric
@@
    Since $\pi:E\to M$ is a fiber bundle associated to the ruling $\mathfrak{S}$, we have $\langle c_{1}(E),[\Sigma_{y}]\rangle=\pm 1$. Therefore, $\pi:\pi^{-1}(\Sigma_{y})\to\Sigma_{y}$ is a circle bundle over the 2-sphere with Euler class $\pm 1$. In other words, it is the Hopf fibration. In particular, $\pi^{-1}(\Sigma_{y})\cong S^{3}$ is simply connected. This proves the inclusion map of the fiber is $\pi_{1}$-null.
@@ 4.1. Corrugation of enlargeable manifolds
The manifold $Q=S^{2}(1/2)\times S^{d-2}(1)$ admits a $1$-small pre-ruling by the spheres $S^{2}(1/2)\times\{*\}$. Some of these spheres intersect the ball $B_{\delta/2}(q)\subset Q$. We isotope the spheres away from this ball by a small isotopy, to obtain a $(1+O(\delta))$-small pre-ruling on $Q\setminus B_{\delta/2}(q)$. Next, for any point $z\in N_{n}\setminus\bigcup_{x\in S_{n}}B_{\delta/2}(x)$, we consider a minimal geodesic ray $\gamma$ from $z$ to the nearest boundary component of $N_{n}\setminus\bigcup_{x\in S_{n}}B_{\delta/2}(x)$. Since $S_{n}$ is $\delta$-dense, length of $\gamma_{z}$ is at most $\delta$. Suppose the terminal point of $\gamma$ is contained in the ruling sphere $S^{2}\cong\Sigma\subset Q\setminus B_{\delta/2}(q)$. Then $\Sigma\cup\gamma$ is a ‘sphere with a hair’ containing $z$. Consider a very thin normal neighborhood $N(\gamma)\cong[0,1]\times D^{d-1}(\delta/100)$ of $\gamma\cong[0,1]$, with $\{0\}\times D^{d-1}$ containing $z$ and $\{1\}\times D^{d-1}\subset\Sigma$. Let $\Sigma_{z}=\partial N(\gamma)~\triangle~\Sigma$ be the symmetric difference. Then $\Sigma_{z}\subset M_{n}$ is an embedded sphere containing $z$. Effectively, $\Sigma_{z}$ is obtained from $\Sigma\cup\gamma$ by replacing the ‘hair’ by a thin ‘finger’ protruding from $\Sigma$. Note that $\Sigma_{z}$ is homologous to $\Sigma\subset Q$, and $\mathrm{diam}(\Sigma_{z})=1+O(\delta)<3/2$.
@@ 5. Main Theorem
Let $\pi:E\to M$ be the circle bundle associated to $\mathfrak{S}$. Then, given any $\sigma>0$, there is a metric $g_{E}$ on $E$, such that $(E,g_{E})$ has $\mathrm{mscal}\geq\sigma$ and $\mathrm{UW}_{d-2}(E)\gtrsim R$.
@@
Next, suppose $\{i,j,k\}\subset\{1,2,3,4\}$. This can be treated in two separate cases. First, suppose $i=1$, $j=2$ and $k\in\{3,4\}$. Then,
@@
Hence, $\langle\psi_{0},J\rangle=1$ in this case. Finally, suppose $i\in\{1,2\}$, $j=3$, $k=4$. Then,

### v2

@@ 1.1. Background and motivation
To remedy this, Alpert-Balitskiy-Guth [ABG24] propose that one should impose a control on the volume of unit balls in the universal cover of a slightly larger concentric ball. We reproduce their proposal as a definition below. Note that the following definition slightly differs from the notion of macroscopic scalar curvature defined in [GUT10a, Section 7].
@@ 1.2. Main results
Certainly, a uniruling is an example of a ruling in this sense. Since $\omega\in H^{2}(M;\mathbf{Z})$, there is a principal circle bundle $\pi:E\to M$ naturally associated with $\omega$, satisfying $c_{1}(E)=\omega$. Furthermore, since $\langle\omega,[\Sigma_{x}]\rangle=\pm 1$, the restriction of $E$ over any ruling sphere is the Hopf fibration. We call $E$ the circle bundle associated to the ruling. Our second key result in this article, used to ensure $E_{n}$ has uniformly positive macroscopic scalar curvature, is the following.
@@
Let $M$ be a closed Riemannian manifold that admits a $3/2$-small ruling. For any $\sigma>0$, there exists a small enough fiber radius $\varepsilon>0$ such that the total space $E$ of the circle bundle $S^{1}({\varepsilon})\to E\to M$ associated with the ruling, equipped with the adapted metric (cf. Definition 2.4) satisfies $\mathrm{mscal}\geq\sigma$.
From Theorem B and Theorem C, we see that Theorem A follows, provided the existence of Riemannian manifolds $M_{n}^{d-1}$ satisfying the following conditions:
@@
    Let $N^{d-1}$ be any $\mathbf{Z}_{2}$-enlargeable (cf. Definition 4.1), $\mathrm{Pin}^{-}$ manifold of dimension $d-1\geq 3$. Then, there exists a cover $N_{n}\to N$ such that $\mathrm{HS}(N_{n};\mathbf{Z}_{2})\geq n$. Choose a uniform small constant $\delta>0$. Let $M_{n}$ be obtained from $N_{n}$ by taking a connect sum with $S^{2}\times S^{d-3}$ at each point of a maximal $\delta$-separated collection of points on $N_{n}$. Then $M_{n}$ admits a (highly singular) small ruling by smoothly embedded $2$-spheres of diameters equal to, upto an error of the order of $\delta$, the diameter of the round $2$-spheres $S^{2}\times\{*\}$ that rule $S^{2}\times S^{d-3}$, which we may choose to be less than $3/2$. The class $\omega\in H^{2}(M_{n};\mathbf{Z})$ may be chosen to be the Poincare dual of the linear combination of the homology classes $\{*\}\times S^{d-3}]$, one from each $S^{2}\times S^{d-3}$ connect summand. For details, see Section 4.1.
@@ Acknowledgments
The authors are grateful to Alexey Balitskiy and Larry Guth for suggesting several ideas and patiently answering their questions at various stages of this work. They are also indebted to Bernhard Hanke and Shihang He for helpful correspondence that led to several key ideas in this article. In particular, He pointed out the article [BOL06] to us, which served as the inspiration for Section 2. They are also grateful to Mike Miller Eismeier for help with the algebraic topology in Section 2, as well as for suggesting several improvements on an earlier draft. They are also thankful to Dennis Sullivan for sharing his insights on symplectic geometry, some of which are behind Section 1.3.
@@ 2.3. Hopf invariant for circle bundles
Let $M^{d}$ be a $d$-dimensional manifold. By Wu’s formula, $\mathrm{Sq}^{2}:H^{d-2}(M;\mathbf{Z}_{2})\to H^{d}(M;\mathbf{Z}_{2})$ is given by $\mathrm{Sq}^{2}(x)=v_{2}(M)\smile x$. Thus, if $M$ admits a $\mathrm{Pin}^{-}$ structure, then $\mathrm{Sq}^{2}=0$ on $H^{d-2}(M;\mathbf{Z}_{2})$.
@@
$\Phi:H^{i}(M;\mathbf{Z}_{2})\to H^{i+2}(\mathrm{Th}(E);\mathbf{Z}_{2})\cong{H^{i+2}(\mathbb{D}(E),\mathbb{S}(E);\mathbf{Z}_{2})},$
@@ 2.4. Width of circle bundles
The results in the following section are not used in the rest of the article. However, it provides an instructive contrast to the results discussed above.
@@ 2.5. Urysohn width and Cheeger-Gromov collapsing
We now fix $n\gg C$. Let $g_{E,\varepsilon}=\pi^{*}g_{Z}+\varepsilon^{2}\theta\otimes\theta$ be the adapted metric on $E$ associated with the same auxiliary connection as $g_{E}$, but with fiber radius $\varepsilon<1$. Let $g_{n,\varepsilon}:=p_{n}^{*}g_{E,\varepsilon}$. Then $g_{n,\varepsilon}<{g_{n}}$, therefore $\mathrm{UW}_{3}(E_{n},g_{n,\varepsilon})<\mathrm{UW}(E_{n},g_{n})\leq C$ as well. Thus, even though there is a Cheeger-Gromov collapse $(E_{n},g_{n,\varepsilon})\xrightarrow{\mathrm{GH}}Z_{n}$, we still have
@@ 3. Singular ruling and macroscopic scalar curvature
Let $(M^{d},g_{M})$ be a closed Riemannian manifold with a $3/2$-small ruling $\mathfrak{S}$. Let $\pi:E\to M$ be the circle bundle associated with $\mathfrak{S}$. We equip $E$ with an adapted metric
@@
    Since $\pi:E\to M$ is a fiber bundle associated with the ruling $\mathfrak{S}$, we have $\langle c_{1}(E),[\Sigma_{y}]\rangle=\pm 1$. Therefore, $\pi:\pi^{-1}(\Sigma_{y})\to\Sigma_{y}$ is a circle bundle over the 2-sphere with Euler class $\pm 1$. In other words, it is the Hopf fibration. In particular, $\pi^{-1}(\Sigma_{y})\cong S^{3}$ is simply connected. This proves the inclusion map of the fiber is $\pi_{1}$-null.
@@ 4.1. Corrugation of enlargeable manifolds
The manifold $Q=S^{2}(1/2)\times S^{d-2}(1)$ admits a $1$-small pre-ruling by the spheres $S^{2}(1/2)\times\{*\}$. Some of these spheres intersect the ball $B_{\delta/2}(q)\subset Q$. We isotope the spheres away from this ball by a small isotopy, to obtain a $(1+O(\delta))$-small pre-ruling on $Q\setminus B_{\delta/2}(q)$. Next, for any point $z\in N_{n}\setminus\bigcup_{x\in S_{n}}B_{\delta/2}(x)$, we consider a minimal geodesic ray $\gamma$ from $z$ to the nearest boundary component of $N_{n}\setminus\bigcup_{x\in S_{n}}B_{\delta/2}(x)$. Since $S_{n}$ is $\delta$-dense, length of $\gamma_{z}$ is at most $\delta$. Suppose the terminal point of $\gamma$ is contained in the ruling sphere $S^{2}\cong\Sigma\subset Q\setminus B_{\delta/2}(q)$. Then $\Sigma\cup\gamma$ is a ‘sphere with a hair’ containing $z$. Consider a very thin normal thickening $N(\gamma)\cong[0,1]\times D^{2}(\delta/100)$ of $\gamma\cong[0,1]$, with $\{0\}\times D^{2}$ containing $z$ and $\{1\}\times D^{2}\subset\Sigma$. Let $\Sigma_{z}=\partial N(\gamma)~\triangle~\Sigma$ be the symmetric difference. Then $\Sigma_{z}\subset M_{n}$ is an embedded sphere containing $z$. Effectively, $\Sigma_{z}$ is obtained from $\Sigma\cup\gamma$ by replacing the ‘hair’ by a thin ‘finger’ protruding from $\Sigma$. Note that $\Sigma_{z}$ is homologous to $\Sigma\subset Q$, and $\mathrm{diam}(\Sigma_{z})=1+O(\delta)<3/2$.
@@ 5. Main Theorem
Let $\pi:E\to M$ be the circle bundle associated with $\mathfrak{S}$. Then, given any $\sigma>0$, there is a metric $g_{E}$ on $E$, such that $(E,g_{E})$ has $\mathrm{mscal}\geq\sigma$ and $\mathrm{UW}_{d-2}(E)\gtrsim R$.
@@
Now, suppose $\{i,j,k\}\subset\{1,2,3,4\}$. Consider the case $i=1$, $j=2$ and $k\in\{3,4\}$. Then,
@@
Hence, $\langle\psi_{0},J\rangle=1$ in this case. Next, consider the case $i\in\{1,2\}$, $j=3$, $k=4$. Then,